what is the answer to 8x-5=7x+7
Answer by bonster(299)

8x-5=7x+7
(8x-5)-7x=(7x+7)-7x subtract 7x from both sides
x-5=7 add 5 to both sides
(x-5)+5=(7)+5
x=12